HUBBLE SPACE TELESCOPE DAILY REPORT #1006 PERIOD COVERED: 0800 EST Daily Status Report as of 306/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make a total of seven internal flat calibrations images. All seven images were reported to be normal. 1.2 Completed FOC/96 5033 (FOC Camera, 2nd, 3rd Stage Focus - Cycle 3 ) The Faint Object Camera (f/96) was used to run the HISTOGRAM mode test of the Camera internal focus. All test exposures were entirely internal using the green LED. This test mode does not produce data that is sent to the ground in the science stream. No anomalies were reported. 1.3 Completed four Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make four sets of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled. A loss of lock occurred during one of the observations following a recentering event. ( See Operational Note 2.1). 1.4 Completed FOS/BL 4122 (The Stellar Content of Wolf - Rayet Galaxies (FOS Spectra)) The Faint Object Spectrograph (Blue Side) was used to make observations of the Wolf - Rayet Galaxy NGC-4214 which is similar to a starbursting galaxy, but has broad emission lines of He II at 4664 A0. Seven observations were received and displayed. The target acquisition was successful. No anomalies were reported. 1.5 Completed PC 4854 (Galaxies and Clusters, WFPC GTO Augmentation, Cycle 3 High) The Planetary Camera was used to make images of the nuclear regions of galaxy NGC-2681-NU to study its luminosity profile. Two observations were monitored in the OSS. The science data still remains on the science tape recorder. No problems were reported. 1.6 Completed HSP/VIS 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make four calibration observations of standard star BD+75D325 to determine the instruments long term sensitivity of the IDT detectors. All four observations were received and displayed. All four observations were reported to look normal. A recentering event occurred during the last observation. 1.7 Completed FOS/RD 4581 (Quasar Absorption Line Survey - Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make observations of quasar PKS2128-12 to learn more about the origin and evolution of gas clouds along the line of sight of this target. The OSS monitored the observation, and assessed the data quality. No anomalies were reported. 1.8 Completed PC 4851 (Globular Cluster Cores, WFPC GTO Augmentation, Cycle 3 High) The Planetary Camera was used to make ultraviolet observations of the core of standard star BD+28D4211 to obtain the point spread function to be used for subsequent globular cluster observations. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 8 Successful : 8 Loss of Fine Lock at 306/0509Z following a recentering event. Subsequent acquisition was successful. 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Note: The GHRS Repair Kit aliveness test was held as scheduled at 6:15PM EST on Monday, Nov. 1. The test sent and verified three lines of GHRS science data to the NSSC-I, using the inflight GHRS instruments and SI C&DH. The test was successful, verifying the aliveness test to be used after the Repair Kit is installed during the First Servicing Mission. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of PC 4644 (The Nuclei of Nearby SO and Spiral Galaxies - I A PC Snapshot Imaging) 3.2 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.3 Complete FOC /96 4800 (The Stellar Content of Wolf - Rayet Galaxies (FOC Images)) 3.4 Complete Two Sets of PC 4714 (WF/PC Cycle 2 Calibration: PC Internal Delta Flats) 3.5 Complete HSP/UV2 1093 (Observations of ZZ Ceti Stars) 3.6 Complete Two Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.7 Complete PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High) 3.8 Complete FOS/RD 3448 (Ultraviolet Evolution of Elliptical Galaxies at Moderate Redshift) 3.9 Complete WFC/PC 5026 (WF/PC Cycle Calibration: Rapid Internal Monitor) 3.10 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.11 Complete FOS/BL 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3: Cycle 3 High) 3.12 Complete HSP/POL 4888 (Remnant Stars in Supernova Remnants - Augmentation - Cycle 3 High) 3.13 Complete PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) 3.14 Complete GHRS/2 2537 (Physical Conditions of Interstellar Gas Within 50 Parsecs) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-